OpenStudy (anonymous):

demonstrate: 1) lim x -> 0 + sqrt x = 0 2) lim x → 0 f (x) = L ⇔ lim x → 0 (f (x) - l) = 0

OpenStudy (anonymous):

Are we going to use the epsilon delta definition?

OpenStudy (anonymous):

The limit: \[ \lim_{x\to a}f(x) = L \]is defined as: \[ \forall \epsilon\;\exists\delta\quad\forall x:\;0\lt |x-a|\lt\delta\implies |f(x)-L|\lt \epsilon \]

OpenStudy (anonymous):

I need help as not prove either

OpenStudy (anonymous):

So you don't need to prove?

OpenStudy (anonymous):

The only way I know to 'demonstrate' it is using Epsilon Delta.

OpenStudy (anonymous):

Not as demonstrate these problems and are for tomorrow

OpenStudy (anonymous):

not as... as not.... what am I supposed to do?

OpenStudy (anonymous):

if you could help me I'd appreciate it

OpenStudy (anonymous):

How do you want help?

OpenStudy (anonymous):

if you can you do the shows to see how you do

OpenStudy (anonymous):

The right-hand limit: \[ \lim_{x\to a^+}f(x) = L \]is defined as: \[ \forall \epsilon\;\exists\delta\quad\forall x:\;0\lt x-a\lt\delta\implies |f(x)-L|\lt \epsilon \] In this case \[ \lim_{x\to 0^+}\sqrt{x} = 0 \]is: \[ \forall \epsilon\;\exists\delta\quad\forall x:\;0\lt x-0\lt\delta\implies |\sqrt{x}-0|\lt \epsilon \]

OpenStudy (anonymous):

We simplify a bit: \[ \forall \epsilon\;\exists\delta\quad\forall x:\;0\lt x\lt\delta\implies |\sqrt{x}|\lt \epsilon \]Since \(x>0\) we can modify the epsilon inequality: \[ \forall \epsilon\;\exists\delta\quad\forall x:\;0\lt x\lt\delta\implies |\sqrt{x}||\sqrt{x}|\lt \epsilon^2 \]\[ \forall \epsilon\;\exists\delta\quad\forall x:\;0\lt x\lt\delta\implies x\lt \epsilon^2 \]

OpenStudy (anonymous):

Notice how the inequality holds when \(\delta = \epsilon^2\)

OpenStudy (anonymous):

So when they give us an \(\epsilon > 0\) we give them \(\delta =\sqrt{\epsilon}\)

OpenStudy (anonymous):

This demonstrates\[ \lim_{x\to 0^+}\sqrt{x} = 0 \]

OpenStudy (anonymous):

Now to demonstrate:\[ \lim_{x \to 0} f (x) = L \iff \lim_{x \to 0} (f (x) - l) = 0 \]

OpenStudy (anonymous):

\[ \lim_{x \to 0} f (x) = L \iff \lim_{x \to 0} (f (x) - L) = 0 \]

OpenStudy (anonymous):

First let's say that \(g(x) = f(x)-L\). \[ \lim_{x \to 0} f (x) = L \iff \lim_{x \to 0} g(x) = 0 \]

OpenStudy (anonymous):

\[ \lim_{x \to 0} f (x) = L \]By definition is: \[ \epsilon>0\;\exists\delta>0\quad\forall x:\;0\lt |x-0|\lt\delta\implies |f(x)-L|\lt \epsilon \]Which is the same as: \[ \epsilon>0\;\exists\delta>0\quad\forall x:\;0\lt |x-0|\lt\delta\implies |g(x)|\lt \epsilon \]Which is the same as\[ \epsilon>0\;\exists\delta>0\quad\forall x:\;0\lt |x-0|\lt\delta\implies |g(x)-0|\lt \epsilon \]Which gets us back to \[ \lim_{x \to 0} g(x) = 0 \]

OpenStudy (anonymous):

Thank you thank you